OpenStudy (anonymous):

Which is an equation of the line passing through points at (1, 6) and (-2, 1)?

OpenStudy (amistre64):

depends on what form they are presenting you with

OpenStudy (amistre64):

at any rate; subtract one point from the other and stack y/x to get the slope

OpenStudy (amistre64):

what would (1, 6) -(-2, 1) ------- equals?

OpenStudy (anonymous):

Slope-int form.

OpenStudy (anonymous):

(-1, 5)?

OpenStudy (amistre64):

close :) i get 3,5 which stack to 5/3 for a slope

OpenStudy (amistre64):

how many options do you have as : y = 5/3 x + b ?

OpenStudy (anonymous):

y = mx + n is the slope intercept form m is the slope m = (y2-y1)/(x2-x1) x1=1, y1=6 x2=-2,y2=1 m = (1-6)/(-2-1) m=-5/-3 m=-5/3

OpenStudy (anonymous):

One option[: Thankyou, haha.

OpenStudy (dumbcow):

slope formula \[m = \frac{y2-y1}{x2-x1}\]

OpenStudy (amistre64):

then that narrows it down :)

OpenStudy (amistre64):

to continue for a dbl chk tho: y = mx -mPx + Py ; where P(x,y) is either point given; ill go with (1, 6) y = (5/3)x -5(1)/3 + 6(3)/3 y = (5/3)x -5/3 + 18/3 y = (5/3)x + 13/3 as long as I did the math right

OpenStudy (anonymous):

You're right![:

OpenStudy (amistre64):

giorgia did fine; but mighta typoed the slope at the end ;)